OpenStudy (anonymous):

Write an equation for the line tangent to the circle x^2 + y^2 = 17, at the point (4, 1). Enter your answer in slope-intercept form: y = mx + b; example: y = -5x-2

OpenStudy (zzr0ck3r):

so you want to take the derivative of the equation of the circle and then that will give the slope \(x^2+y^2=17\implies 2x+2y\frac{dy}{dx}=0\\\frac{dy}{dx}=\frac{-x}{y}\) so at the point \((4,1)\) the slope is \(-4\) so your line will be of the form \[\large y=-4x+b\] now use the point \((4,1)\) to find the value of b note we know that the tangent line touches the circle at \((4,1)\) so we know that point is also on our line.

OpenStudy (anonymous):

So 17?

OpenStudy (anonymous):

for the value of B ?

OpenStudy (anonymous):

@zzr0ck3r :o

OpenStudy (zzr0ck3r):

how did you get 17?

OpenStudy (anonymous):

Idk :/ I plugged in 4 for x and 1 for y

OpenStudy (zzr0ck3r):

yes you are right im sorry its early lol

OpenStudy (anonymous):

Omg don't be sorry thanks for starting me off, I was struggling with this. thank you so much:)